solve A 1.2 kw immersion heater is use to supply energy for 75 seconds. the energy supplied is use to heat 800g of paraffin in calorimeter of heat capacity 42.5 j/k from 20 degree Celsius to 66 degree Celsius. calculate the specific heat capacity of the paraffin


1
Expert's answer
2021-05-19T11:05:36-0400
"Pt=C\Delta T+mc_p\Delta T\\(1200)(75)=42.5(66-20)+c_p(66-20)(0.8)\\c_p=2390\frac{J}{kgK}"